# A Variant of the Bravyi-Terhal Bound for Arbitrary Boundary Conditions

### v1

@@ 1.1 Motivations
In recent years, quantum Low-Density Parity-Check (qLDPC) codes have become of great interest since they are a promising ingredient for scalable quantum computing. While random matrices are commonly used to build classical LDPC codes, due to an orthogonality constraint, an algebraic ansatz is required to build quantum LDPC codes, which makes the task challenging. The last few years have seen spectacular breakthroughs, with Panteleev and Kalachev proving the existence of asymptotically good quantum LDPC codes [25], see also [26]. However, it is not clear whether these constructions can give the best codes for short, practical lengths. For short lengths, significant efforts have gone into the search for good Two-Block codes, as is apparent from [31] in particular.
These quantum binary LDPC codes have parity-check matrices constructed from a pair $(\mathbf{A},\mathbf{B})$ of square commuting matrices. The first family of such codes was originally introduced by MacKay et al. [11]. These quantum codes are obtained by taking $\mathbf{B}=\mathbf{A}^{\intercal}$ a cyclic matrix, and were named Bicycle codes. An interesting feature of this construction is that the resulting quantum CSS codes are obtained from classical self-dual codes given by the parity-check matrices $[\mathbf{A},\mathbf{A}^{\intercal}]$. A generalised version of this construction was presented by Pryadko and Kovalev [19]. These Generalised Bicycle codes are obtained by taking $\mathbf{A}$ and $\mathbf{B}$ both cyclic. For the sake of simplicity, we will from now on also refer to these codes as Bicycles codes. A broader generalisation was presented by Pryadko and Lin [30] in the form of Two-Block Group Algebra (2BGA) codes, constructed from a group algebra of a finite group.
Some of these quantum codes were extensively studied in [19, 22, 27] and demonstrate promising error-correcting capabilities. In the best cases, they offer almost the same error rate threshold as the surface codes, whilst improving the dimension and encoding overhead [31]. However, the behaviour of the minimum distance of this family of codes is unknown and remains an intriguing problem.
Abelian Two-Block Group Algebra codes generalise Kitaev’s toric codes [8, 6, 15], since the latter correspond to the case where the abelian group is chosen to be a product of two cyclic groups. Moreover, Abelian Two-Block Group Algebra codes have the potential to yield codes with even larger minimum distances than the Kitaev code. Additionally, like toric codes, these codes can similarly be represented on lattices, albeit of higher dimensions, in such a way that the stabilizer generators of the code only act on a fixed number of qubits, all present in local regions. This feature may be appealing for physical implementations but also imply limitations on the code parameters, which are the subject of the present paper.
@@ 1.2 Previous results
Due to physical constraints, a lot of attention has gone into stabilizer codes whose generators are local in a strongly geometric sense. These codes are designed so that each stabilizer generator affects only a limited number of nearby qubits. Bravyi and Terhal [16] studied the constraints on the parameters of these codes, providing a crucial insight into the limitations of stabilizer codes. They proved that for stabilizer codes whose qubits are mapped to a region of $\mathbb{Z}^{D}$ such as $\{1,\ldots,L\}^{D}$, or with a periodic structure such as $(\mathbb{Z}/L\mathbb{Z})^{D}$, where each stabilizer generator’s support is included within a cube of $\rho^{D}$ vertices, then, whenever $L\geq 2(\rho-1)^{2}$, the minimum distance of the code cannot exceed
@@
Subsequent research, such as [28, 29], extended this analysis to quantum LDPC codes without explicit lattice embeddings, but with more general graph embeddings for which the subexponential behaviour of the minimum distance can also be shown, but with loosely specified constants.
Appealing to the original Bravyi-Terhal bound, Pryadko and Wang [27] gave an upper bound on the minimum distances of Bicycle codes, a subclass of Abelian 2BGA codes. It states that the minimum distance of non-trivial Bicycle codes of length $N$ and stabilizer generators of weight $w$ are in $O(N^{\frac{D-1}{D}})$ where $D\leq w-1$ and $D\leq w-2$ in some cases, with unspecified constants.
@@ 1.3 Contributions
In the present paper, we bridge the gap between the previously known results [16, 28, 29] on geometrically-local quantum codes and revisit the Bravyi-Terhal bound for quantum LDPC codes that can be defined by local checks on a quotient of a $D$-dimensional lattice. More precisely, consider a quantum LDPC code of length $N=nm$. If the physical qubits of the code can be indexed by the vertices of the quotient of a $D$-dimensional lattice in such a way that each vertex indexes $m$ qubits and such that all the stabilizer generators are supported within a ball of radius $\rho$, then the minimum distance of the code satisfies $d\leq m\sqrt{\gamma_{D}}(\sqrt{D}+4\rho)n^{\frac{D-1}{D}}$ whenever $n^{1/D}\geq 8\rho\sqrt{\gamma_{D}}$, where $\gamma_{D}$ is the $D$-dimensional Hermite constant. See Theorem 3.6 for the formal statement.
@@
As an application, we provide upper bounds on the minimum distance of all Abelian 2BGA codes. Precisely, applying Theorem 3.6 with $m=2$ and $\rho=1$, we show that any Abelian 2BGA code with code length $2n$ and stabilizer generators of weight $w$, the minimum distance is bounded from above by $2\sqrt{\gamma_{D}}(\sqrt{D}+4)n^{\frac{D-1}{D}}$ whenever $n^{1/D}\geq 8\sqrt{\gamma_{D}}$, where $D=w-2$ and $\gamma_{D}$ is the $D$-dimensional Hermite constant. This bound is more explicit and somewhat sharper than the minimum distance bound given in [27] for Bicycle codes: it also applies to the more general class of Abelian 2BGA codes.
@@ 2.1 Stabilizer and CSS codes
A quantum Calderbank-Shor-Steane (CSS) code [3, 4] is a stabilizer code for which each generator is a product of only bit-flip operators $X$ or a product of only phase-flip operators $Z$. It follows that its parity-check matrix is of the form
@@
The length $N$ of a CSS code, or the number of physical qubits, is the number of columns of $\mathbf{H}_{X}$ (or of $\mathbf{H}_{Z}$). Its dimension $k$, or the number of logical qubits, is given by $k=N-\mathrm{rank}\,\mathbf{H}_{X}-\mathrm{rank}\,\mathbf{H}_{Z}$. The minimum distance $d$ of the code is defined as $\min(d_{X},d_{Z})$, where $d_{X}$ (resp. $d_{Z}$) is equal to the smallest Hamming weight of a non-zero vector of $\ker\mathbf{H}_{X}$ (resp. $\ker\mathbf{H}_{Z})$ that is not in the row-space of $\mathbf{H}_{Z}$ (resp. $\mathbf{H}_{X})$.
@@ 2.2 Abelian 2BGA codes
Inspired by the previously known Bicycles codes [19], Abelian Two-Block Group Algebra (2BGA) codes were proposed by Priadko and Lin [30]. These quantum binary LDPC codes are defined by parity-check matrices constructed from a pair of square commuting matrices obtained from a group algebra of a finite abelian group.
@@
An Abelian 2BGA code is a CSS stabilizer code defined by the two binary matrices $\mathbf{H}_{X}=[\mathbf{A}|\mathbf{B}]$, $\mathbf{H}_{Z}=[\mathbf{B}^{\intercal}|\mathbf{A}^{\intercal}]$ where $\mathbf{A}=\sum_{g\in G}a_{g}\mathbb{B}(g)$ and $\mathbf{B}=\sum_{g\in G}b_{g}\mathbb{B}(g)$, with $a_{g},b_{g}\in\mathbb{F}_{2}$. Such a code is non-trivial whenever $\mathbf{A}$ or $\mathbf{B}$ is non-zero. The code length is $N=2n$. Clearly $\mathrm{rank}\,\mathbf{H}_{X}=\mathrm{rank}\,\mathbf{H}_{Z}$ and so the code dimension $k$ is even [21]. Moreover, we have $d_{X}=d_{Z}$ since the Abelian 2BGA codes defined by the matrices $(\mathbf{A},\mathbf{B})$ and $(\mathbf{B},\mathbf{A})$ are equivalent [30].
When the abelian group is taken to be cyclic, one obtains the Bicycle codes introduced by Pryadko and Kovalev in [19]. If furthermore $\mathbf{A}=\mathbf{B}$, then one recovers the Bicycle Codes of MacKay et al. [11]. When the abelian group is the product of two cyclic groups, i.e. of the form $\mathbb{Z}_{r}\times\mathbb{Z}_{s}$, then one obtains the Bivariate Bicycle codes [31]. Finally, note that Abelian 2BGA codes are the simplest form of Lifted Product codes introduced by Panteleev and Kalachev [24], with both matrices of dimension $1\times 1$.
@@ 3.2 Hermite and Rankin constants
Definition 3.3 (Hermite constant [1])
@@
Rankin [2] provided a generalisation of Hermite constants.
@@
The Rankin constants are defined by $\gamma_{D,r}=\max\{\gamma_{D,r}(\Lambda)\,|\,\Lambda\text{ rank $D$ lattice}\}$. By a duality argument [7], it can be shown that $\gamma_{D,r}=\gamma_{D,D-r}$. Moreover, note that $\gamma_{D,1}$ is clearly equal to $\gamma_{D}$, the $D$-dimensional Hermite constant.
@@ 3.3 Geometrically-local codes
A stabilizer code is said to be geometrically-local if one can index its physical qubits by the vertices of a lattice in such a way that the code’s stabilizer generators act only on neighbouring qubits. In [16], Bravyi and Terhal showed that when such a geometrically-local code can be mapped to a region of $\mathbb{Z}^{D}$ such as $\{1,\ldots,L\}^{D}$, or with a periodic structure such as $(\mathbb{Z}/L\mathbb{Z})^{D}$, an upper bound on the minimum distance of the code is given by (1). We consider a variation of their theorem, for which the code is mapped on a quotient $\mathbb{Z}^{D}/\Lambda$, where $\Lambda$ is an arbitrary $D$-dimensional sublattice of $\mathbb{Z}^{D}$.
@@
Let us start by recalling the Cleaning lemma that was first introduced by Bravyi and Terhal in [16]. Let $\mathcal{C}$ be a stabilizer code with stabilizer group $S$. Let $M$ denote a subset of the $N$ qubits in the code block. If $\mathcal{O}$ is one of the code’s logical operators, we say that $\mathcal{O}$ can be cleaned on $M$ if there is a logically equivalent operator $\tilde{\mathcal{O}}=\mathcal{O}\mathcal{S}$, where $\mathcal{S}$ is an element of the code stabilizer $S$, such that $\tilde{\mathcal{O}}$ acts trivially on $M$.
@@
To prove Theorem 3.6, we shall use the Cleaning Lemma repeatedly. We therefore need to define appropriate regions that partition the qubits of the code $\mathcal{C}$. The first step of the proof is to choose an appropriate basis for our lattice. In our case, it is a basis whose associated Gram-Schmidt basis has a vector of large enough length. A convenient way to find such basis is to use the work of [17] in order to build a family $\{u_{1},\dots,u_{D-1}\}$ of linearly independent vectors of $\Lambda$ that can be completed into a basis of $\Lambda$ and which is such that the volume of the $D-1$ dimensional lattice generated by $u_{1},\dots,u_{D-1}$ is minimal.
@@
Proof: From the definition of the Rankin constant [7], we have that $\gamma_{D,D-1}(\Lambda)\leq\gamma_{D}$ and that there is a rank $(D-1)$ sublattice $S_{0}\subset\Lambda$ such that $0<\gamma_{D,D-1}(\Lambda)=\mathrm{vol}(S)^{2}\,n^{-2\,\frac{D-1}{D}}$. We now need to find a basis $u_{1},\dots,u_{D}$ of $\Lambda$ such that $\mathrm{vol}(S_{0})=\mathrm{vol}(\{u_{1},\dots,u_{D-1}\})$. We prove the existence of this basis by using the structure theorem for finitely generated modules over the principal ideal domain $\mathbb{Z}$. Since $\Lambda$ and $S_{0}$ are respectively lattices of rank $D$ and $D-1$, there exist a basis $\{u_{1},\dots,u_{D}\}$ of $\Lambda$ and integers $\0x_{1},\dots,x_{D-1}\in\mathbb{Z}\backslash\{0\}$ such that $\{x_{1}u_{1},\dots,x_{D-1}u_{D-1}\}$ is a basis of $S_{0}$. We have that the volume of $S_{0}$ is equal to $\mathrm{vol}(\{u_{1},\dots,u_{D-1}\})$. Indeed, let $T$ be the $D-1$ lattice generated by $\{u_{1},\dots,u_{D-1}\}$. On one hand, by minimality of the volume of $S_{0}$, we have that $\mathrm{vol}(S_{0})\leq\mathrm{vol}(T)$. On the other hand, since $S_{0}$ is generated by $\{x_{1}u_{1},\dots,x_{D-1}u_{D-1}\}$, we have $\mathrm{vol}(S_{0})=|T/S_{0}|\mathrm{vol}(T)\geq\mathrm{vol}(T)$. It follows that $\mathrm{vol}(S_{0})=\mathrm{vol}(T)=\mathrm{vol}(\{u_{1},\dots,u_{D-1}\})$. . $\square$
Let $u_{1},\dots,u_{D}\in\Lambda$ be a basis of $\Lambda$ such that $\Lambda$ such that $Vol(u_{1},\dots,u_{D-1})=\sqrt{\gamma_{D,D-1}(\Lambda)}n^{\frac{D-1}{D}}$. This particular basis has the property that the length of the last vector of the associated orthogonal Gram-Schmidt basis $\{u^{\star}_{1},\ldots,u^{\star}_{D}\}$ is as large as possible.
@@
$\Gamma=\{\,\sum_{i=1}^{D}\,a_{i}u_{i}\,\,|\,\,\forall i\in[[1,D-1]],0\leq a_{i}<1\,\text{ and }\,\frac{-\sqrt{D}}{2\|u_{D}^{*}\|}+\frac{k}{\mu}\leq a_{D}<\frac{\sqrt{D}}{2\|u_{D}^{*}\|}+\frac{k+1}{\mu}\,.\}$
@@
$\mathrm{vol}(\Gamma/\Lambda)$ $=\mathrm{vol}(\,\mathcal{P}\,\cap\,\cup_{g\in\Lambda}(g+\Gamma)\,)$
$\leq\sum_{g\in\Lambda}\mathrm{vol}(\mathcal{P}\cap\Gamma+g)=\mathrm{vol}(\Gamma)$
@@
Proof: If $n^{1/D}\geq 8\rho\sqrt{\gamma_{D}}$, then from Lemma 3.9, we have $\ell\geq\frac{n^{1/D}}{\sqrt{\gamma_{D}}}\geq 8\rho$. Hence $\mu\geq\lfloor\frac{\ell}{2\rho}\rfloor-1>\frac{\ell}{2\rho}-2\geq 2$. It follows that $\lambda=\frac{\ell}{\mu}<\frac{2\rho\ell}{\ell-4\rho}\leq 4\rho$ since $0<2\rho\ell\leq 4\rho(\ell-4\rho)$. . $\square$

### v2

@@ 1.1 Motivations
In recent years, quantum Low-Density Parity-Check (qLDPC) codes have become of great interest since they are a promising ingredient for scalable quantum computing. While random matrices are commonly used to build classical LDPC codes, due to an orthogonality constraint, an algebraic ansatz is required to build quantum LDPC codes, which makes the task challenging. The last few years have seen spectacular breakthroughs, with Panteleev and Kalachev proving the existence of asymptotically good quantum LDPC codes [PK22a], see also [LZ22]. However, it is not clear whether these constructions can give the best codes for short, practical lengths. For short lengths, significant efforts have gone into the search for good Two-Block codes, as is apparent from [BCG+24] in particular.
These quantum binary LDPC codes have parity-check matrices constructed from a pair $(\mathbf{A},\mathbf{B})$ of square commuting matrices. The first family of such codes was originally introduced by MacKay et al. [MMM04]. These quantum codes are obtained by taking $\mathbf{B}=\mathbf{A}^{\intercal}$ a cyclic matrix, and were named Bicycle codes. An interesting feature of this construction is that the resulting quantum CSS codes are obtained from classical self-dual codes given by the parity-check matrices $[\mathbf{A},\mathbf{A}^{\intercal}]$. A generalised version of this construction was presented by Pryadko and Kovalev [KP13]. These Generalised Bicycle codes are obtained by taking $\mathbf{A}$ and $\mathbf{B}$ both cyclic. For the sake of simplicity, we will from now on also refer to these codes as Bicycles codes. A broader generalisation was presented by Pryadko and Lin [LP24] in the form of Two-Block Group Algebra (2BGA) codes, constructed from a group algebra of a finite group.
Some of these quantum codes were extensively studied in [KP13, PK21, WP22] and demonstrate promising error-correcting capabilities. In the best cases, they offer almost the same error rate threshold as the surface codes, whilst improving the dimension and encoding overhead [BCG+24]. However, the behaviour of the minimum distance of this family of codes is unknown and remains an intriguing problem.
Abelian Two-Block Group Algebra codes generalise Kitaev’s toric codes [Kit03, DKLP02, BMD07], since the latter correspond to the case where the abelian group is chosen to be a product of two cyclic groups. Moreover, Abelian Two-Block Group Algebra codes have the potential to yield codes with even larger minimum distances than the Kitaev code. Additionally, like toric codes, these codes can similarly be represented on lattices, albeit of higher dimensions, in such a way that the stabilizer generators of the code only act on a fixed number of qubits, all present in local regions. This feature may be appealing for physical implementations but also imply limitations on the code parameters, which are the subject of the present paper.
@@ 1.2 Previous results
Due to physical constraints, a lot of attention has gone into stabilizer codes whose generators are local in a strongly geometric sense. These codes are designed so that each stabilizer generator affects only a limited number of nearby qubits. Bravyi and Terhal [BT09] studied the constraints on the parameters of these codes, providing a crucial insight into the limitations of stabilizer codes. They proved that for stabilizer codes whose qubits are mapped to a region of $\mathbb{Z}^{D}$ such as $\{1,\ldots,L\}^{D}$, or with a periodic structure such as $(\mathbb{Z}/L\mathbb{Z})^{D}$, where each stabilizer generator’s support is included within a cube of $\rho^{D}$ vertices, then, whenever $L\geq 2(\rho-1)^{2}$, the minimum distance of the code cannot exceed
@@
Subsequent research, such as [BK22, BGKL24], extended this analysis to quantum LDPC codes without explicit lattice embeddings, but with more general graph embeddings for which the subexponential behaviour of the minimum distance can also be shown, but with loosely specified constants.
Appealing to the original Bravyi-Terhal bound, Pryadko and Wang [WP22] gave an upper bound on the minimum distances of Bicycle codes, a subclass of Abelian 2BGA codes. It states that the minimum distance of non-trivial Bicycle codes of length $N$ and stabilizer generators of weight $w$ are in $O(N^{\frac{D-1}{D}})$ where $D\leq w-1$ and $D\leq w-2$ in some cases, with unspecified constants.
@@ 1.3 Contributions
In the present paper, we bridge the gap between the previously known results [BT09, BK22, BGKL24] on geometrically-local quantum codes and revisit the Bravyi-Terhal bound for quantum LDPC codes that can be defined by local checks on a quotient of a $D$-dimensional lattice. More precisely, consider a quantum LDPC code of length $N=nm$. If the physical qubits of the code can be indexed by the vertices of the quotient of a $D$-dimensional lattice in such a way that each vertex indexes $m$ qubits and such that all the stabilizer generators are supported within a ball of radius $\rho$, then the minimum distance of the code satisfies $d\leq m\sqrt{\gamma_{D}}(\sqrt{D}+4\rho)n^{\frac{D-1}{D}}$ whenever $n^{1/D}\geq 8\rho\sqrt{\gamma_{D}}$, where $\gamma_{D}$ is the $D$-dimensional Hermite constant. See Theorem 3.6 for the formal statement.
@@
As an application, we provide upper bounds on the minimum distance of all Abelian 2BGA codes. Precisely, applying Theorem 3.6 with $m=2$ and $\rho=1$, we show that any Abelian 2BGA code with code length $2n$ and stabilizer generators of weight $w$, the minimum distance is bounded from above by $2\sqrt{\gamma_{D}}(\sqrt{D}+4)n^{\frac{D-1}{D}}$ whenever $n^{1/D}\geq 8\sqrt{\gamma_{D}}$, where $D=w-2$ and $\gamma_{D}$ is the $D$-dimensional Hermite constant. This bound is more explicit and somewhat sharper than the minimum distance bound given in [WP22] for Bicycle codes: it also applies to the more general class of Abelian 2BGA codes.
@@ 2.1 Stabilizer and CSS codes
A quantum Calderbank-Shor-Steane (CSS) code [CS96, Ste96] is a stabilizer code for which each generator is a product of only bit-flip operators $X$ or a product of only phase-flip operators $Z$. It follows that its parity-check matrix is of the form
@@
The length $N$ of a CSS code, or the number of physical qubits, is the number of columns of $\mathbf{H}_{X}$ (or of $\mathbf{H}_{Z}$). Its dimension $k$, or the number of logical qubits, is $k=N-\mathrm{rank}\,\mathbf{H}_{X}-\mathrm{rank}\,\mathbf{H}_{Z}$. The minimum distance $d$ of the code is defined as $\min(d_{X},d_{Z})$, where $d_{X}$ (resp. $d_{Z}$) is equal to the smallest Hamming weight of a non-zero vector of $\ker\mathbf{H}_{X}$ (resp. $\ker\mathbf{H}_{Z})$ that is not in the row-space of $\mathbf{H}_{Z}$ (resp. $\mathbf{H}_{X})$.
@@ 2.2 Abelian 2BGA codes
Inspired by the previously known Bicycles codes [KP13], Abelian Two-Block Group Algebra (2BGA) codes were proposed by Priadko and Lin [LP24]. These quantum binary LDPC codes are defined by parity-check matrices constructed from a pair of square commuting matrices obtained from a group algebra of a finite abelian group.
@@
An Abelian 2BGA code is a CSS stabilizer code defined by two binary matrices $\mathbf{H}_{X}=[\mathbf{A}|\mathbf{B}]$, $\mathbf{H}_{Z}=[\mathbf{B}^{\intercal}|\mathbf{A}^{\intercal}]$ where $\mathbf{A}=\sum_{g\in G}a_{g}\mathbb{B}(g)$ and $\mathbf{B}=\sum_{g\in G}b_{g}\mathbb{B}(g)$, with $a_{g},b_{g}\in\mathbb{F}_{2}$. Such a code is non-trivial whenever $\mathbf{A}$ or $\mathbf{B}$ is non-zero. The code length is $N=2n$. Clearly $\mathrm{rank}\,\mathbf{H}_{X}=\mathrm{rank}\,\mathbf{H}_{Z}$ and so the code dimension $k$ is even [PK20]. Moreover, we have $d_{X}=d_{Z}$ since the Abelian 2BGA codes defined by the matrices $(\mathbf{A},\mathbf{B})$ and $(\mathbf{B},\mathbf{A})$ are equivalent [LP24].
When the abelian group is taken to be cyclic, one obtains the Bicycle codes introduced by Pryadko and Kovalev in [KP13]. If furthermore $\mathbf{A}=\mathbf{B}$, then one recovers the Bicycle Codes of MacKay et al. [MMM04]. When the abelian group is the product of two cyclic groups, i.e. of the form $\mathbb{Z}_{r}\times\mathbb{Z}_{s}$, then one obtains the Bivariate Bicycle codes [BCG+24]. Finally, note that Abelian 2BGA codes are the simplest form of Lifted Product codes introduced by Panteleev and Kalachev [PK22b], with both matrices of dimension $1\times 1$.
@@ 3.2 Hermite and Rankin constants
Definition 3.3 (Hermite constant [Her50])
@@
Rankin [Ran56] provided a generalisation of Hermite constants.
@@
The Rankin constants are defined by $\gamma_{D,r}=\max\{\gamma_{D,r}(\Lambda)\,|\,\Lambda\text{ rank $D$ lattice}\}$. By a duality argument [Mar02], it can be shown that $\gamma_{D,r}=\gamma_{D,D-r}$. Moreover, note that $\gamma_{D,1}$ is clearly equal to $\gamma_{D}$, the $D$-dimensional Hermite constant.
@@ 3.3 Geometrically-local codes
A stabilizer code is said to be geometrically-local if one can index its physical qubits by the vertices of a lattice in such a way that the code’s stabilizer generators act only on neighbouring qubits. In [BT09], Bravyi and Terhal showed that when such a geometrically-local code can be mapped to a region of $\mathbb{Z}^{D}$ such as $\{1,\ldots,L\}^{D}$, or with a periodic structure such as $(\mathbb{Z}/L\mathbb{Z})^{D}$, an upper bound on the minimum distance of the code is given by (1). We consider a variation of their theorem, for which the code is mapped on a quotient $\mathbb{Z}^{D}/\Lambda$, where $\Lambda$ is an arbitrary $D$-dimensional sublattice of $\mathbb{Z}^{D}$.
@@
Let us start by recalling the Cleaning lemma that was first introduced by Bravyi and Terhal in [BT09]. Let $\mathcal{C}$ be a stabilizer code with stabilizer group $S$. Let $M$ denote a subset of the $N$ qubits in the code block. If $\mathcal{O}$ is one of the code’s logical operators, we say that $\mathcal{O}$ can be cleaned on $M$ if there is a logically equivalent operator $\tilde{\mathcal{O}}=\mathcal{O}\mathcal{S}$, where $\mathcal{S}$ is an element of the code stabilizer $S$, such that $\tilde{\mathcal{O}}$ acts trivially on $M$.
@@
To prove Theorem 3.6, we shall use the Cleaning Lemma repeatedly. We therefore need to define appropriate regions that partition the qubits of the code $\mathcal{C}$. The first step of the proof is to choose an appropriate basis for our lattice. In our case, it is a basis whose associated Gram-Schmidt basis has a vector of large enough length. A convenient way to find such basis is to use the work of [Ngu09] in order to build a family $\{u_{1},\dots,u_{D-1}\}$ of linearly independent vectors of $\Lambda$ that can be completed into a basis of $\Lambda$ and which is such that the volume of the $D-1$ dimensional lattice generated by $u_{1},\dots,u_{D-1}$ is minimal.
@@
Proof: From the definition of the Rankin constant [Mar02], we have that $\gamma_{D,D-1}(\Lambda)\leq\gamma_{D}$ and that there is a rank $(D-1)$ sublattice $S_{0}\subset\Lambda$ such that $0<\gamma_{D,D-1}(\Lambda)=\mathrm{vol}(S)^{2}\,n^{-2\,\frac{D-1}{D}}$. We now need to find a basis $u_{1},\dots,u_{D}$ of $\Lambda$ such that $\mathrm{vol}(S_{0})=\mathrm{vol}(\{u_{1},\dots,u_{D-1}\})$. We prove the existence of this basis by using the structure theorem for finitely generated modules over the principal ideal domain $\mathbb{Z}$. Since $\Lambda$ and $S_{0}$ are respectively lattices of rank $D$ and $D-1$, there exist a basis $\{u_{1},\dots,u_{D}\}$ of $\Lambda$ and integers $\0x_{1},\dots,x_{D-1}\in\mathbb{Z}\backslash\{0\}$ such that $\{x_{1}u_{1},\dots,x_{D-1}u_{D-1}\}$ is a basis of $S_{0}$. We have that the volume of $S_{0}$ is equal to $\mathrm{vol}(\{u_{1},\dots,u_{D-1}\})$. Indeed, let $T$ be the $D-1$ lattice generated by $\{u_{1},\dots,u_{D-1}\}$. On one hand, by minimality of the volume of $S_{0}$, we have that $\mathrm{vol}(S_{0})\leq\mathrm{vol}(T)$. On the other hand, since $S_{0}$ is generated by $\{x_{1}u_{1},\dots,x_{D-1}u_{D-1}\}$, we have $\mathrm{vol}(S_{0})=|T/S_{0}|\mathrm{vol}(T)\geq\mathrm{vol}(T)$. It therefore follows that $\mathrm{vol}(S_{0})=\mathrm{vol}(T)=\mathrm{vol}(\{u_{1},\dots,u_{D-1}\})$. . $\square$
Let $u_{1},\dots,u_{D}\in\Lambda$ be a basis of $\Lambda$ such that $\mathrm{vol}(u_{1},\dots,u_{D-1})=\sqrt{\gamma_{D,D-1}(\Lambda)}n^{\frac{D-1}{D}}$. This particular basis has the property that the length of the last vector of the associated orthogonal Gram-Schmidt basis $\{u^{\star}_{1},\ldots,u^{\star}_{D}\}$ is as large as possible.
@@
$\Gamma=\{\,\sum_{i=1}^{D}\,a_{i}u_{i}\,\,|\,\,\forall i\in[[1,D-1]],0\leq a_{i}<1\,\text{ and }\,\frac{-\sqrt{D}}{2\|u_{D}^{*}\|}+\frac{k}{\mu}\leq a_{D}<\frac{\sqrt{D}}{2\|u_{D}^{*}\|}+\frac{k+1}{\mu}\}.$
@@
$\mathrm{vol}(\Gamma/\Lambda)=\mathrm{vol}(\,\mathcal{P}\,\cap\,\cup_{g\in\Lambda}(g+\Gamma)\,)\leq\sum_{g\in\Lambda}\mathrm{vol}(\mathcal{P}\cap\Gamma+g)=\mathrm{vol}(\Gamma).$
@@
Proof: If $n^{1/D}\geq 8\rho\sqrt{\gamma_{D}}$, then from Lemma 3.9, we have $\ell\geq\frac{n^{1/D}}{\sqrt{\gamma_{D}}}\geq 8\rho$. Hence we obtain $\mu\geq\lfloor\frac{\ell}{2\rho}\rfloor-1>\frac{\ell}{2\rho}-2\geq 2$. It follows that $\lambda=\frac{\ell}{\mu}<\frac{2\rho\ell}{\ell-4\rho}\leq 4\rho$ since $0<2\rho\ell\leq 4\rho(\ell-4\rho)$. . $\square$